# MINT: Multi-Vector Search Index Tuning

### v1

@@ 5.2. End-to-end Performance
The end-to-end latency, speedup, and cost are in Figure 7 (using HNSW) and Figure 8 (using DiskANN). Mint and PerColumn use a fixed storage budget that equals the number of columns. PerQuery, as a reference, uses additional storage (33% more on Naïve and 50% more on BiSimple, BiComplex, and News). Across all workloads and indexes, Mint consistently outperforms PerColumn in terms of latency. The speedup ranges from 2.1×2.1$\times$2.1 × to 7.2×7.2$\times$7.2 × for HNSW indexes and 2.1×2.1$\times$2.1 × to 8.3×8.3$\times$8.3 × for DiskANN indexes. In addition, Mint achieves comparable latency compared to the latency lower bound obtained from PerQuery in almost all scenarios. Mint outperforms PerColumn because Mint has a lower cost in each workload. Due to space limitations, we show each query’s latency in the BiComplex workload as an example in Figure 9. Mint outperforms PerColumn on all queries using HNSW and DiskANN. The speedup ranges from 1.5×1.5$\times$1.5 × to 27.8×27.8$\times$27.8 × with median=11.1×11.$1$.
The performance gains of Mint stem from its use of multi-column indexes. For example, in Figure 9B, PerColumn has much higher latency than Mint, especially on queries with more columns, such as Query 7 (6 columns, 138 ms), Query 2 (4 columns, 76 ms), and Query 3 (3 columns, 106 ms). In comparison, Mint uses 4-column and 3-column indexes to answer these three queries, achieving latency speedups from 6.8×6.8$\times$6.8 × to 27.8×27.8$\times$27.8 ×. In addition, these multi-column indexes are shared by multiple queries, which helps Mint meet the storage constraint.
Figure 8. Using DiskANN, (A) Mint outperforms PerColumn on all four workloads, and Mint’s latency is close to the optimal PerQuery’s, although PerQuery uses 33% to 50% more storage. (B) The speedup ranges from 2.1×2.1$\times$2.1 × to 8.3×8.3$\times$8.3 ×. (C) Mint outperforms PerColumn because Mint has lower cost.
Figure 9. Query latency for (A) HNSW and (B) DiskANN in BiComplex workload and (C) queries probabilities.
@@
Figure 10. Mint finds configurations with lower latency when storage increases.
@@
We present a case study to show how Mint benefits from multi-column plans. Table 3 shows the query plans for the Naïve workload using DiskANN. The columns GloVe100, SIFT1M, and Yandex Text-to-Image are numbered $1,2$, and $3$. Each plan shows the indexes followed by the numbers of extracted items (i.e., $ek$). Each query’s latency is in Figure 11.
@@
Figure 11. Query latency in Naïve workload using DiskANN.
@@
Table 3. Compared to single-column plans, multi-column plans extract much fewer items to meet recall constraint in Naïve workload.
5.3. Scalability Analysis
In this section, we show how the runtime of Mint is affected by the storage constraint and number of queries.
@@ 5.3. Scalability Analysis
As shown in Figure 14, increasing the storage budget has a minimal impact on the total planning time. Most intermediate results are generated in the early iterations and are cached using the mechanism in Section 4.2. Thus, they can be reused for subsequent iterations. This allows more indexes to be created, increasing the storage budget without substantially increasing computation costs.
@@
Figure 14 shows that the running time of Mint grows almost linearly with the number of queries in a workload. This is because the runtime of Algorithm 3 is dominated by getting plans for each query given the current set of indexes (Line 3), and the runtime of this step grows linearly with the number of queries.
@@
Figure 12. Mint finds configurations in minutes. The running time breaks down into one-time training and index tuning. The one-time training runs only once given a database, and its time can be amortized across different workloads.
@@
Figure 13. Mint’s running time does not increase much when storage increases, because storage is within a limited range in practice.
@@
Figure 14. As workload size increases, the running time increases almost linearly.
{mdframed}
@@
Key Takeaway. Mint is efficient and scales well with storage constraints and workload sizes.
6. Related Work
@@
The single-vector search is known as Nearest Neighbor search and dates back to 1960s (Cover and Hart, 2006). Researchers have proposed various techniques including $k$-d tree (Bentley, 1975), ball tree (Omohundro, 1989), etc. Due the growth of vector sizes for many scenarios, such as RAG, exact nearest neighbor search is infeasible. Therefore, researchers have developed Approximate Nearest Neighbor (ANN) search techniques to retrieve nearest neighbors more efficiently.
There are several categories of ANN techniques (Pan et al., 2024a). Partition-based techniques (Chen et al., 2021; Babenko and Lempitsky, 2014; Lv et al., 2007; Liu et al., 2014; Norouzi et al., 2012; Datar et al., 2004) put similar items into the same partition. Given a query, they limit the scan within a few promising partitions, e.g., those whose centroids are closest to the query. Partitions may or may not overlap. For instance, partitions using $k$-means clustering (Lloyd, 1982) can be non-overlapping, and partitions using hashing (Andoni and Indyk, 2008) can be overlapping.
Tree-based techniques (Muja and Lowe, 2009; Silpa-Anan and Hartley, 2008; Dasgupta and Freund, 2008) split items recursively into a tree. The split strategy can be based on axis, plane, etc. Graph-based techniques are prevalent these days because of their efficiency and high recall (Aumüller et al., 2020; Simhadri et al., 2022, 2024). Examples include DiskANN (Subramanya et al., 2019), HNSW (Malkov and Yashunin, 2018), NSW (Malkov et al., 2014), and NSG (Fu et al., 2017). Quantization techniques (Gray, 1984; Jegou et al., 2010; Ge et al., 2013; Guo et al., 2020) can be used by all these methods to compress vectors and improve efficiency. They are implemented in various industry libraries and systems (Chen et al., 2018; Douze et al., 2024; Johnson et al., 2019; Boytsov and Naidan, 2013; Yang et al., 2020; Guo et al., 2022; Su et al., 2024; PgV, [n.d.]; Pin, [n.d.]; Zil, [n.d.]; Luc, [n.d.]; Azu, [n.d.]).
All of the above ANN techniques focus on single-vector search. They are orthogonal to our Multi-Vector Search Index Tuning problem and can be plugged into our flexible framework.
@@
In multi-vector search, items are represented by multiple vectors (Pan et al., 2024a; Wang et al., 2024; Duong et al., 2021), each of which is in a separate column. The one-index-per-column approach creates an index on each column and scan multiple indexes to answer a query. For instance, Milvus (Wang et al., 2021b) and VBase (Zhang et al., 2023) scan multiple single-column indexes in turn and merge the results to find top results. OneSparse (Chen et al., 2024) pushes down an intersection operation across indexes to filter out low-quality results earlier. Compared to scanning single-column indexes, we scan multi-column indexes to find the top result more quickly. In addition, we study which multi-column indexes should be created and scanned across multiple queries.
@@
Fagin et al.’s threshold algorithms (Fagin et al., 2001, 2003) and follow-on work (Badr and Vodislav, 2015; Theobald et al., 2004; Chen et al., 2011) address the problem of answering exact and approximate top-k queries. They scan multiple sorted lists in turns until the result cannot be improved by scanning the remaining items. These query processing algorithms are orthogonal to our Mint algorithm, which focuses on index tuning. If these algorithms can provide cost and recall estimators, then they can be plugged into our framework.
@@
The Index Tuning problem has been extensively studied for relational databases (Wu et al., 2024; Chaudhuri and Narasayya, 2007). Examples include AutoAdmin (Chaudhuri and Narasayya, 1997), CoPhy (Dash et al., 2011), and Relaxation (Bruno and Chaudhuri, 2005). This problem significantly differs from multi-vector search index tuning in its cost model and query planning. Therefore, we needed to develop new algorithms for multi-vector search index tuning.
7. Conclusions
@@
In this paper, we propose Mint, a framework for multi-vector search index tuning problem. In a multi-vector database where each rows is an item, each column represents a feature, and each cell is a high-dimensional vector. Given a workload of queries, Mint recommends indexes to minimize the latency while meeting the storage and recall constraints. Mint consists of two major modules, Query Planner and Configuration Searcher. We formally define the problems of Query Planning and Configuration Searching, prove they are NP-hard, and develop algorithms to solve them. Our experiments show a 2.1×2.1$\times$2.1 × to 8.3×8.3$\times$8.3 × latency speedup compared to the state of the art. Since Mint is designed to be flexible with estimators and ANN indexes, in future work, we plan to develop more accurate estimators for more performance gain and apply Mint to other types of indexes in addition to HNSW and DiskANN.

### v2

@@ V-B End-to-end Performance
[…]
The performance gains of Mint stem from its use of multi-column indexes. For example, in Figure 9B, PerColumn has much higher latency than Mint, especially on queries with more columns, such as Query 7 (6 columns, 138 ms), Query 2 (4 columns, 76 ms), and Query 3 (3 columns, 106 ms). In comparison, Mint uses 4-column and 3-column indexes to answer these three queries, achieving latency speedups from $6.8\times$ to $27.8\times$. In addition, these multi-column indexes are shared by multiple queries, which helps Mint meet the storage constraint.
Figure 8: Using DiskANN, (A) Mint outperforms PerColumn on all four workloads, and Mint’s latency is close to the optimal PerQuery’s, although PerQuery uses 33% to 50% more storage. (B) The speedup ranges from $2.1\times$ to $8.3\times$. (C) Mint outperforms PerColumn because Mint has lower cost.
Figure 9: Query latency for (A) HNSW and (B) DiskANN in BiComplex workload and (C) queries probabilities.
@@
Figure 10: Mint finds configurations with lower latency when storage increases.
@@
We present a case study to show how Mint benefits from multi-column plans. Table III shows the query plans for the Naïve workload using DiskANN. The columns are numbered $1,2$, and $3$. Each plan shows the indexes followed by the numbers of extracted items (i.e., $ek$). Each query’s latency is in Figure 11.
@@
Figure 11: Query latency in Naïve workload using DiskANN.
@@
TABLE III: Compared to single-column plans, multi-column plans extract much fewer items to meet recall constraint in Naïve workload.
V-C Scalability Analysis
Here we show how the runtime of Mint is affected by the storage constraint and number of queries, and show that Mint still outperforms PerColumn on a large-scale dataset.
@@ V-C Scalability Analysis
As shown in Figure 13, increasing the storage budget has a minimal impact on the total planning time. Most intermediate results are generated in the early iterations and are cached using the mechanism in Section IV-B. Thus, they can be reused for subsequent iterations. This allows more indexes to be created, increasing the storage budget without substantially increasing computation costs.
@@
Figure 14 shows that the running time of Mint grows almost linearly with the number of queries in a workload. This is because the runtime of Algorithm 2 is dominated by getting plans for each query given the current set of indexes (Line 2), and the runtime of this step grows linearly with the number of queries.
@@
Figure 12: Mint finds configurations in minutes. The running time breaks down into one-time training and index tuning. The one-time training runs only once given a database, and its time can be amortized across different workloads.
@@
Figure 13: Mint’s running time does not increase much when storage increases, because storage is within a limited range in practice.
@@
Figure 14: As workload size increases, the running time increases almost linearly.
Large-scale Dataset
[…]
@@
Key Takeaway. Mint is efficient and scales well with storage constraints and workload sizes, and still outperforms PerColumn on 10M rows.
VI Related Work
@@
The single-vector search is known as Nearest Neighbor search and dates back to 1960s [23, 10, 62]. Due to the growth of vector sizes for many scenarios, such as RAG, exact nearest neighbor search is infeasible. Therefore, researchers have developed Approximate Nearest Neighbor (ANN) search techniques to retrieve nearest neighbors more efficiently.
There are several categories of ANN techniques [63]. Partition-based techniques [21, 7, 54, 52, 60, 26, 53, 2] put similar items into the same partition. Tree-based techniques [59, 71, 24] split items recursively into a tree. Graph-based techniques organize items in graphs and are prevalent these days because of their efficiency and high recall [5, 73, 72, 44, 55, 56, 33]. Quantization techniques [37, 45, 36, 39, 35, 34] can be used by all these methods to compress vectors and improve efficiency. They are implemented in various industry libraries and systems [20, 28, 48, 12, 84, 38, 74, 66, 67, 90, 3, 6]. All of the above ANN techniques focus on single-vector search. They are orthogonal to our Multi-Vector Search Index Tuning problem and can be plugged into our flexible framework.
@@
In multi-vector search, items are represented by multiple vectors [63, 78, 29, 27], each of which is in a separate column. The one-index-per-column approach creates an index on each column and scans multiple indexes to answer a query. For instance, Milvus [77] and VBase [86] scan multiple single-column indexes in turn and merge the results to find top results. OneSparse [22] pushes down an intersection operation across indexes to filter out low-quality results earlier. Compared to scanning single-column indexes, we scan multi-column indexes to find the top result more quickly. In addition, we study which multi-column indexes should be created and scanned across multiple queries.
Another line of work represents each item as a set of multiple vectors within a single semantic domain (e.g., token- or paragraph-level embeddings) and computes similarity using late-interaction functions such as MaxSim [49, 69, 70, 11]. Each item consists of a single column whose value is a variable-size set of vectors. These methods focus on fine-grained matching over a single column and do not consider multiple heterogeneous vector columns or the index tuning problem studied in this work, where queries may involve different subsets of columns.
Some related works study multi-column vector queries but do not consider index tuning. DEG [85] focuses on weighted queries over exactly two fixed vector columns. HJG [89] proposes a well-designed unified index that supports weighted queries over multiple vector columns, but it relies on a single fixed index and does not consider workload-aware index tuning. In practice, queries may have heterogeneous recall requirements and benefit from dedicated indexes on specific column subsets, which can achieve lower latency and better recall than a unified index. Our work instead studies index tuning for a query workload, selecting and configuring indexes under varying recall constraints.
@@
Fagin et al.’s threshold algorithms [30, 31] and follow-on work [8, 76, 19] address the problem of answering exact and approximate top-k queries. They scan multiple sorted lists in turns until the result cannot be improved by scanning the remaining items. These query processing algorithms are orthogonal to our Mint algorithm, which focuses on index tuning. If these algorithms can provide cost and recall estimators, then they can be plugged into our framework.
@@
The Index Tuning problem has been extensively studied for relational databases [82, 18, 16, 25, 13]. This problem significantly differs from multi-vector search index tuning in its cost model and query planning. Therefore, we needed to develop new algorithms for multi-vector search index tuning.
An index can be viewed as a special case of a materialized view (MV) [1, 40]. While MVs are effective for queries involving filters, aggregations, and joins over multiple tables, our work focuses on vector indexes on a single table without such operations; thus, we do not consider MVs as a separate design choice. MVs may become relevant in extended settings such as filtered vector search [51], which we leave as future work.
VII Conclusions
@@
In this paper, we propose Mint, a framework for the multi-vector search index tuning problem. In a multi-vector database, each row is an item, each column represents a feature, and each cell is a high-dimensional vector. Given a workload of queries, Mint recommends indexes to minimize the latency while meeting the storage and recall constraints. Mint has two major modules, Query Planner and Configuration Searcher. We formally define the problems of Query Planning and Configuration Searching, prove they are NP-hard, and develop algorithms to solve them. Our experiments show a $2.1\times$ to $8.3\times$ latency speedup compared to the state of the art. Since Mint is designed to be flexible with estimators and ANN indexes, in future work, we plan to develop more accurate estimators for more performance gain and apply Mint to other types of indexes in addition to HNSW and DiskANN.
VIII AI-Generated Content Acknowledgment
No section in this paper was AI-generated.